OpenStudy (anonymous):

Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. Point B: (3,5) Point C: (5,3)

OpenStudy (anonymous):

I know it's y=mx+b I just need help plugging the numbers in

OpenStudy (anonymous):

would it be 5=1/2+ whatever b is...not sure?

OpenStudy (anonymous):

Use this : m=Y2-Y1/X2-X1 X1=3 Y1=5 We just want (-2,3) Caz formula is this : Y2-Y=m(X2-X) X2=5 Y2=3

OpenStudy (anonymous):

But I thought slope intercept form way y=mx+b?

OpenStudy (anonymous):

*wass

hero (hero):

It is...and you can use it to find the values of \(m\) and \(b\)

OpenStudy (anonymous):

So m is not 1/2 in this one?

hero (hero):

Let's see...plug in the points: 3 = 5m + b 5 = 3m + b b = b 3 - 5m = 5 - 3m 3 - 5 = 5m - 3m -2 = 2m -2/2 = m -1 = m Looks like the slope is m = -1

OpenStudy (anonymous):

oh okay so would y be 5, like 5=-1+ then what ever b is?

hero (hero):

Basically, there is a specific set sequence of steps to use to write the equation of a line in point-slope form. But, in short, all you need to do is find the values of \(m\) and \(b\) then plug them in to the original formula.

hero (hero):

We found \(m\) but we need to find \(b\).

hero (hero):

In order to find \(b\) we will use the value of the slope along with one of the given points and insert them into the formula. Let's just use (3,5): \(y = mx + b\) \(3 = -1(5) + b\)

OpenStudy (anonymous):

so Would it be like 3=5b?

hero (hero):

You are changing the equation

hero (hero):

It's 3 = -5 + b

hero (hero):

You can't just get rid of the plus

hero (hero):

And you can't multiply 5 with b

OpenStudy (anonymous):

So how do I get b If I can't do anything to it?

hero (hero):

You have to get b by itself. It's called "isolating"

hero (hero):

To get b by itself, you have to add 5 to both sides: 3 = -1(5) + b 3 = -5 + b 5 + 3 = b 8 = b

OpenStudy (anonymous):

Oh duh...I feel smart :).. so In the equation would it be 5=-1+8?

OpenStudy (anonymous):

or would 3 be y?

hero (hero):

The equation of a line has to be in the form y = mx + b

hero (hero):

You don't get rid of any of the variables, however, you can replace m and b with their equivalent values.

hero (hero):

Remember, that's what I was telling you all along. That we will find the values of \(m\) and \(b)\) then insert those values in to the formula.

hero (hero):

\(y = \color\red{m}x + \color\red{b}\)

hero (hero):

The variables \(m\) and \(b\) are replace with the values we found for them.

OpenStudy (anonymous):

okay so y=-1x+8

hero (hero):

Yes, finally...

OpenStudy (anonymous):

YEAH! Thank you so much... You've been so helpful, I suck at math lol :)